OpenStudy (anonymous):

Define the following vectoraddition and scalarmultiplication on R^2

OpenStudy (anonymous):

\[(x_1,y_1)+(x_2,y_2)= (x_1+x_2+1,y_1+y_2)\]

OpenStudy (anonymous):

and

OpenStudy (anonymous):

\[a(x,y)=(ax,ay)\]

OpenStudy (amistre64):

\[(x_1,y_1)+(x_2,y_2)= (x_1+x_2+1,y_1+y_2)\] ^^??

OpenStudy (anonymous):

Do these mathematical applications determine a vectorspace on R2?

OpenStudy (anonymous):

the +1 is kinda weird to me

OpenStudy (amistre64):

a vector space must have an indentity, an inverse, and a zero vector; right? and be closed

OpenStudy (anonymous):

an identity?

OpenStudy (amistre64):

a + i = a ; and element that produces no results

OpenStudy (amistre64):

think of 0 for addition ,a dn 1 for multiplication

OpenStudy (anonymous):

oh yes

OpenStudy (amistre64):

(a,b) + (xi, yi) a + xi + 1 = a; xi = -1 b + yi = b; yi = 0

OpenStudy (amistre64):

since the multiplication is defined as usual for a vector space ... there is no mystery about it

OpenStudy (amistre64):

(a,b) + (xz,yz) = 0,0 a + xz + 1 = 0 xz = -(a+1) b + yz = 0; yz = -b

OpenStudy (anonymous):

with i, you just mean a random integer that doesn't change the vector?

OpenStudy (amistre64):

xi is an indentity element ... x sub i, or we could have said ix, its just convenient letters to describe the term employed

OpenStudy (anonymous):

lost me there

OpenStudy (anonymous):

googled it, i get it

OpenStudy (amistre64):

:)

OpenStudy (anonymous):

so how..i don't know how to solve this, how to think of a solution even

OpenStudy (amistre64):

i just demonstrated how to determine identity and inverse. Since the elements to pull from come from the set of real numbers then all these operations are possible; the multiplication stuff is just the usual vector scalar definition so there is no new information to add to that part

OpenStudy (amistre64):

might want to see if commutative and associate work out

OpenStudy (anonymous):

alright

OpenStudy (amistre64):

let x = (xa, xb), let y = (ya,yb) x+y = (xa+ya+1 , xb + yb) does y + x equal the same thing? y+x = (xa+ya+1 , xb + yb) yes for some z = (za,zb) does the associative property hold?

OpenStudy (amistre64):

(x+y)+z = ((xa+ya+1)+za+1 , xb + yb + zb) x+(y+z) = the same?

OpenStudy (amistre64):

y+z = (ya+za+1, yb + zb) x + (y+z) = (xa +1 + (ya+za+1), xb + yb + zb)

OpenStudy (anonymous):

z..?

OpenStudy (anonymous):

to prove asso. ?

OpenStudy (amistre64):

yeah, the associative property requires at least 3 elements to assess if: (x+y)+z = x+(y+z), then the associative property holds

OpenStudy (anonymous):

all this proving is still kinda difficult for me

OpenStudy (anonymous):

i'll write down what you wrote and try and understand

OpenStudy (amistre64):

you just gotta keep the definitions that they use in defining the operations; if you can produce the same "form" each time then you have a closed operation

OpenStudy (anonymous):

hmm ok

OpenStudy (amistre64):

good luck

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

i really don't get it